OpenStudy (anonymous):

Anyone good with Quadratic Formulas? I have a couple of questions for you. x^2+9x-4=0 -4x^2+5x-3=0

OpenStudy (tyteen4a03):

You just throw in the numbers. For an equation ax^2 + bx + c = 0, \(x = \frac{-b \pm \sqrt{b^2-4ac}} {2a}\)

OpenStudy (anonymous):

I really don't understand.

OpenStudy (tyteen4a03):

@Kagome_gurl8 Which part of it confuses you?